# Agnostic Process Tomography

### v1

@@ 1.3 Our Results
$\mathrm{d_{avg}}(\mathcal{E}_{1},\mathcal{E}_{2})\triangleq\sqrt{\mathop{\mathbb{E}}_{\ket{\psi}\sim\mu_{S}}[\mathrm{d_{tr}}(\mathcal{E}_{1}(\outerproduct{\psi}{\psi}),\mathcal{E}_{2}(\outerproduct{\psi}{\psi}))^{2}]},$ (1.3)
where the expectation is taken over Haar-random states, and $\mathrm{d_{tr}}$ denotes the trace distance $\mathrm{d_{tr}}(\rho,\sigma)\triangleq\norm{\rho-\sigma}_{1}/2$. For operators on an $n$-qubit space, the normalized Frobenius distance is
$d_{F}(A,B)\triangleq\frac{1}{\sqrt{2^{n}}}\norm{A-B}_{F}.$ (1.4)
Meanwhile, for superoperators in an $n$-qubit space, the normalized Frobenius distance is
$d_{F}(\Phi,\Psi)\triangleq\frac{1}{2^{n}\sqrt{2}}\norm{J(\Phi)-J(\Psi)}_{F},$ (1.5)
@@
For agnostic unitary estimation, we consider the classes of circuits consisting of $G$ 2-qubit gates [27], Pauli strings, unitary $k$-juntas, and degree-$d$ unitaries. Pauli strings are unitaries of the form $\{I,X,Y,Z\}^{\otimes n}$. Unitary $k$-juntas act nontrivially only on $k$ out of $n$ qubits (Definition 15). Also, degree-$d$ unitaries are unitaries whose Pauli coefficients of order greater than $d$ are 00 (Definition 16).
For agnostic process tomography, we consider similar classes, namely circuits consisting of $G$ gates, Pauli strings, Pauli channels, $k$-junta channels, degree-$d$ channels, and $\mathsf{QAC}^{0}$ channels [22]. Here, Pauli channels are defined as those of the form $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$ for a probability distribution $\{p_{x}\}_{x}$. $k$-junta channels and degree-$d$ channels are defined similarly to the unitary case (see Definitions 18 and 19). Finally, we also consider $n$-to-$1$ $\mathsf{QAC}^{0}$ channels [22], which are defined by applying a unitary implemented by a $\mathsf{QAC}^{0}$ circuit [86] and tracing out all but one qubit (see Definition 21).
Moreover, under compatibility conditions on the concept classes, we show that one can convert any agnostic state tomography algorithm (with respect to fidelity) into an agnostic unitary estimation/process tomography algorithm (with respect to the normalized Frobenius distance $d_{F}$) by learning the Choi state of the unknown unitary/channel. More precisely, if $\mathcal{C}$ is the concept class for agnostic state tomography, this conversion holds when, for each unitary $U$ in the concept class for agnostic unitary estimation/process tomography, the Choi state of $U$ is in the class $\mathcal{C}$. Note that for agnostic process tomography, the concept class must be unitary. This immediately gives us $1$-agnostic unitary estimation algorithms for Clifford circuits, Clifford circuits with few T gates, and circuits consisting of tensor products of single-qubit gates from existing agnostic state tomography algorithms [43] with a quadratic increase with respect to the error $\epsilon$. We obtain similar results for agnostic process tomography with a weaker guarantee on the error ($\mathsf{err}\leq 2^{1/4}\mathsf{opt}^{1/2}+\epsilon$). We note that we can avoid the quadratic increase in $\epsilon$ and obtain 1-agnostic unitary estimation/process tomography with respect to entanglement infidelity, defined as the infidelity between two Choi states (Equation 4.30). This measure has been considered widely in (realizable) unitary estimation [87, 6, 88, 89, 90, 91, 92, 93], but we focus on $d_{F}$ for consistency with our other results.
@@
Degree-$d$ Unitaries (Thm. 3) Improper 1-agnostic $(d_{F})$ $d\cdot 16^{d}n^{2d}\log(n)/\epsilon^{2}$ $\mathcal{O}(N)$
@@
Degree-$d$ Channels (Thm. 8) Improper 1-agnostic ($d_{F}$) $d\cdot 256^{d}n^{2d}\log(n)/\epsilon^{2}$ $\mathcal{O}(N)$
@@
For both agnostic unitary estimation and agnostic process tomography, the algorithms for circuits consisting of $G$ gates are sample-efficient but computationally inefficient. In fact, we prove a more general result (see Propositions 1 and 3): for any finite-sized class $\mathcal{C}$, one can design an agnostic learning algorithm that utilizes $\mathcal{O}(\log|\mathcal{C}|)$ queries to the unknown quantum process but has time complexity scaling with $|\mathcal{C}|$. Qualitatively, this tells us that efficient agnostic tomography is easy from a sample complexity perspective, but it can be difficult to achieve with respect to time complexity. This is similar to an analogous argument for agnostic state tomography [42] based on shadow tomography [94]. Thus, for agnostic process tomography, one should focus on computational efficiency rather than sample efficiency. This is in contrast to the realizable setting, where results typically focus on sample complexity.
@@ 1.4 Proof Ideas
Fourier analysis of Boolean functions [95] is widely used in theoretical computer science and classical learning theory. In particular, it has been used to develop efficient learning algorithms for functions with a sufficiently sparse Fourier decomposition [96, 97, 98] and in classical agnostic learning [47, 99, 100, 101].
Pauli spectrum analysis [28] generalizes Fourier analysis for quantum operators (see Section 2.2 for a more thorough discussion). Analogous to the Fourier decomposition, the key idea is that one can uniquely decompose an operator $A$ acting on an $n$-qubit system into a linear combination of Pauli strings:
@@
where $\sigma_{x}\in\{I,X,Y,Z\}^{\otimes n}$. Recently, [25] further generalized this to the Fourier spectrum of superoperators (see Section 2.3), where a superoperator $\Phi$ in an $n$-qubit space is uniquely decomposed as
@@
where $J(\Phi)$ denotes the Choi represenation of $\Phi$, i.e., the unnormalized Choi state. As in classical learning theory, many quantum learning algorithms utilize Pauli/Fourier spectrum analysis [28, 24, 22, 30, 29, 40, 25].
@@
To resolve this, for agnostic unitary estimation, we “project” our improper hypothesis to its closest unitary using a technique from [23]. We show that this outputs the closest unitary with respect to the normalized Frobenius distance $d_{F}$, resulting in proper agnostic unitary estimation. However, this problem is more difficult for agnostic process tomography. We discuss a potential approach based on convex optimization [102] for projecting an improper hypothesis onto a valid quantum channel [^1] Note that ideas from the quasiprobability decomposition and probabilistic error cancellation which approximate linear superoperators by linear combinations of quantum channels [103, 104, 105, 81, 106, 107, 108, 109, 110] do not help because we need to exactly output a quantum channel. in Section 5.7. We are unable to fully solve the problem due to one difficult-to-verify condition for the convergence guarantees of [102] to hold (Conjecture 1). Nevertheless, we made substantial progress towards resolving this open question, which is also recurrent in the realizable process tomography literature [22, 24, 25]. Now, we summarize our results and the techniques used for each class.
@@
First, we consider any finite-sized class $\mathcal{C}$ of unitaries. Similarly to [27], we use Clifford classical shadows [111] to estimate the average infidelity between the unknown quantum channel $\Phi$ and all $U_{i}\in\mathcal{C}$. Then, we output the unitary in $\mathcal{C}$ with the smallest estimated distance. Thus, we solve proper agnostic unitary estimation for $\mathcal{C}$ with respect to the root mean squared trace distance $\mathrm{d_{avg}}$ and show this uses $\mathcal{O}(\log|\mathcal{C}|)$ queries to $\Phi$ and time $\mathcal{O}(|\mathcal{C}|)$. For many classes of interest, this results in a sample-efficient but computationally inefficient algorithm, emphasizing the importance of computational efficiency for agnostic learning. We show a weaker version of this result for agnostic process tomography of $\mathcal{C}$, which stems from converting from average infidelity to average trace distance for mixed inputs. As a corollary, by utilizing known covering number bounds [27], we obtain agnostic learning algorithms for the class of unitaries implementable by $G$ two-qubit gates. Note that in the realizable setting, [27] proves computational lower bounds. We prove this result in Sections 4.1 and 5.1.
@@
Recently, agnostic state tomography for several concept classes with respect to fidelity [42, 43] has been studied. We show that this translates into guarantees for agnostic unitary estimation/process tomography for a compatible concept class $\mathcal{C}$ of unitaries with respect to $d_{F}$ with a quadratic increase in the error $\epsilon$. In particular, we require that the Choi state of each $U\in\mathcal{C}$ is in the concept class for agnostic state tomography. Our algorithm is simple: apply the agnostic state tomography algorithm to the Choi state of the unknown quantum process. Specifically, we obtain improper $1$-agnostic unitary estimation but a weaker error bound for agnostic process tomography (see Table 1(b)), again due to converting between infidelity and trace distance for mixed inputs. The key step in our proof shows a relationship between the infidelity and the normalized Frobenius distance between two Choi states. We prove this result in Sections 4.2 and 5.2.
@@
Next, we consider the class of Pauli strings $\{I,X,Y,Z\}^{\otimes n}$. We use the fact that one can efficiently sample from the distribution formed by the Pauli/Fourier coefficients [28]. After collecting samples, we simply output the most frequently sampled Pauli string. Thus, we achieve proper 1-agnostic unitary estimation/process tomography. We prove this in Sections 4.3 and 5.3.
@@
Pauli channels are a naturally occurring and widely studied [112, 113, 31, 114, 81, 85, 115, 31, 33, 34, 35, 36, 37, 32] class of quantum channels. Pauli channels are of the form $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$ for a probability distribution $\{p_{x}\}_{x}$. Similarly to Pauli strings, we can sample from $\{p_{x}\}_{x}$. Then, agnostic tomography of Pauli channels reduces to empirically estimating $\{p_{x}\}_{x}$. Thus, we obtain proper 1-agnostic process tomography of Pauli channels. We prove this in Section 5.3.
@@
A $k$-junta unitary/channel is an $n$-qubit unitary/channel that acts only on $k$ qubits (Definitions 15 and 18). [24, 25] devised realizable algorithms for learning junta unitaries/channels. Here, they estimate the influence (Equation 2.16) of each qubit, then select a highly influential subset of qubits and perform full state tomography on the Choi state restricted to these qubits.
@@
[30] develops realizable algorithms leveraging Bohnenblust-Hille (BH) inequalities for operators [116]/
superoperators [30]. These inequalities provide upper bounds on norms of the spectra of low-degree objects. Using these bounds, [30] show that it suffices to learn significantly fewer Pauli/Fourier coefficients, resulting in algorithms with complexities independent of $n$. Using similar arguments for agnostic unitary estimation, we show that the algorithm of [30] is inefficient and only performs better than the trivial algorithm when the optimal error of the class of low-degree unitaries is at most inverse exponential in $n$, i.e., when the problem is almost realizable. We remark that improved bounds for the BH inequality may result in better guarantees for this algorithm and leave this open for future work. As this algorithm does not work well in the simpler unitary case, we do not analyze its performance for agnostic process tomography. We prove these results in Sections 4.5 and 5.5.
@@
The circuit class $\mathsf{QAC}^{0}$ consists of constant-depth circuits with controlled-$Z$ gates and arbitrary single-qubit gates [86]. In [22], the authors study the class of channels formed by applying a $\mathsf{QAC}^{0}$ circuit and tracing out all but one qubit. [22] shows that the Choi representations of such channels have Pauli weight concentrated on low-degree terms. Similarly to [22], we learn the low-degree coefficients of the Choi representation of an arbitrary channel using classical shadows [111], resulting in improper 1-agnostic process tomography for this class of $\mathsf{QAC}^{0}$ channels. We show this in Section 5.6.

### v2

@@ 1.3 Our Results
[…]
where the expectation is taken over Haar-random states, and $\mathrm{d_{tr}}$ denotes the trace distance $\mathrm{d_{tr}}(\rho,\sigma)\triangleq\mathinner{\!\left\lVert\rho-\sigma\right\rVert}_{1}/2$. For operators on an $n$-qubit space, the normalized Frobenius distance is
$d_{F}(A,B)\triangleq\frac{1}{\sqrt{2^{n}}}\mathinner{\!\left\lVert A-B\right\rVert}_{F},$ (1.4)
where $\|.\|_{F}$ is the Frobenius norm (see Section 2.2). Meanwhile, for superoperators in an $n$-qubit space, the normalized Frobenius distance is
$d_{F}(\Phi,\Psi)\triangleq\frac{1}{2^{n}\sqrt{2}}\mathinner{\!\left\lVert J(\Phi)-J(\Psi)\right\rVert}_{F},$ (1.5)
@@
For agnostic unitary estimation, we consider the classes of circuits consisting of $G$ 2-qubit gates [zhao2023learning], Pauli strings, unitary $k$-juntas, and degree-$d$ unitaries. Pauli strings are unitaries of the form $\{I,X,Y,Z\}^{\otimes n}$. Unitary $k$-juntas act nontrivially only on $k$ out of $n$ qubits (Definition 15). Also, degree-$d$ unitaries are unitaries whose Pauli coefficients of order greater than $d$ are 0 (Definition 16).
For agnostic process tomography, we consider similar classes, namely circuits consisting of $G$ gates, Pauli strings, Pauli channels, $k$-junta channels, degree-$d$ channels, and $\mathsf{QAC}^{0}$ channels [nadimpalli2024pauli]. Here, Pauli channels are defined as those of the form $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$ for a probability distribution $\{p_{x}\}_{x}$. $k$-junta channels and degree-$d$ channels are defined similarly to the unitary case (see Definitions 18 and 19). Finally, we also consider $n$-to-11 $\mathsf{QAC}^{0}$ channels [nadimpalli2024pauli], which are defined by applying a unitary implemented by a $\mathsf{QAC}^{0}$ circuit [moore1999quantum] and tracing out all but one qubit (see Definition 21).
Moreover, under compatibility conditions on the concept classes, we show that one can convert any agnostic state tomography algorithm (with respect to fidelity) into an agnostic unitary estimation/process tomography algorithm (with respect to the normalized Frobenius distance $d_{F}$) by learning the Choi state of the unknown unitary/channel. More precisely, if $\mathcal{C}$ is the concept class for agnostic state tomography, this conversion holds when, for each unitary $U$ in the concept class for agnostic unitary estimation/process tomography, the Choi state of $U$ is in the class $\mathcal{C}$. Note that for agnostic process tomography, the concept class must be unitary. This immediately gives us $1$-agnostic unitary estimation algorithms for Clifford circuits, Clifford circuits with few T gates, and circuits consisting of tensor products of single-qubit gates from existing agnostic state tomography algorithms [chen2024stabilizer] with a quadratic increase with respect to the error $\epsilon$. We obtain similar results for agnostic process tomography with a weaker guarantee on the error ($\mathsf{err}\leq 2^{1/4}\mathsf{opt}^{1/2}+\epsilon$). We note that we can avoid the quadratic increase in $\epsilon$ and obtain 1-agnostic unitary estimation/process tomography with respect to entanglement infidelity, defined as the infidelity between two Choi states (Equation 4.30). This measure has been considered widely in (realizable) unitary estimation [leung2000towards, haah2023query, acin2001optimal, peres2002covariant, hayashi2006parallel, chiribella2005optimal, kahn2007fast, yang2020optimal], but we focus on $d_{F}$ for consistency with our other results.
@@
Degree-$d$ Unitaries (Thm. 3) Improper 1-agnostic $(d_{F})$ $\mathcal{O}\left(d\cdot 9^{d}n^{2d}\log(n)/\epsilon^{2}\right)$ $\mathcal{O}(N)$
@@
Degree-$d$ Channels (Thm. 8) Improper 1-agnostic ($d_{F}$) $\mathcal{O}\left(d\cdot 81^{d}n^{4d}\log(n)/\epsilon^{2}\right)$ $\mathcal{O}(N)$
@@
For both agnostic unitary estimation and agnostic process tomography, the algorithms for circuits consisting of $G$ gates are sample-efficient but computationally inefficient. In fact, we prove a more general result (see Propositions 1 and 3): for any finite-sized class $\mathcal{C}$, one can design an agnostic learning algorithm that utilizes $\mathcal{O}(\log|\mathcal{C}|)$ queries to the unknown quantum process but has time complexity scaling with $|\mathcal{C}|$. Qualitatively, this tells us that efficient agnostic tomography is easy from a sample complexity perspective, but it can be difficult to achieve with respect to time complexity. This is similar to an analogous argument for agnostic state tomography [grewal2024agnostic] based on shadow tomography [aaronson2018shadow]. Thus, for agnostic process tomography, one should focus on computational efficiency rather than sample efficiency. This is in contrast to the realizable setting, where results typically focus on sample complexity.
@@ 1.4 Proof Ideas
Fourier analysis of Boolean functions [o2014analysis] is widely used in theoretical computer science and classical learning theory. In particular, it has been used to develop efficient learning algorithms for functions with a sufficiently sparse Fourier decomposition [goldreich1989hard, kushilevitz1991learning, linial1993constant] and in classical agnostic learning [kearns1992toward, heidari2023agnostic, goldwasser2021interactive, gur2024power].
Pauli spectrum analysis [montanaro2008quantum] generalizes Fourier analysis for quantum operators (see Section 2.2 for a more thorough discussion). Analogous to the Fourier decomposition, the key idea is that one can uniquely decompose an operator $A$ acting on an $n$-qubit system into a linear combination of Pauli strings:
@@
where $\sigma_{x}\in\{I,X,Y,Z\}^{\otimes n}$. Recently, [bao2023testing] further generalized this to the Fourier spectrum of superoperators (see Section 2.3), where a superoperator $\Phi$ in an $n$-qubit space is uniquely decomposed as
@@
where $J(\Phi)$ denotes the Choi represenation of $\Phi$, i.e., the unnormalized Choi state. As in classical learning theory, many quantum learning algorithms utilize Pauli/Fourier spectrum analysis [montanaro2008quantum, chen2023testing, nadimpalli2024pauli, arunachalam2024learning, rouze2024quantum, chen2024predicting, bao2023testing].
@@
To resolve this, for agnostic unitary estimation, we “project” our improper hypothesis to its closest unitary using a technique from [huang2024learning]. We show that this outputs the closest unitary with respect to the normalized Frobenius distance $d_{F}$, resulting in proper agnostic unitary estimation. However, this problem is more difficult for agnostic process tomography. We discuss a potential approach based on convex optimization [wang2024convergence] for projecting an improper hypothesis onto a valid quantum channel [^1] Note that ideas from the quasiprobability decomposition and probabilistic error cancellation which approximate linear superoperators by linear combinations of quantum channels [jiang2021physical, regula2021operational, temme2017error, van2023probabilistic, zhao2023power, horodecki2002method, horodecki2003limits, fiuravsek2002structural, korbicz2008structural] do not help because we need to exactly output a quantum channel. in Section 5.7. We are unable to fully solve the problem due to one difficult-to-verify condition for the convergence guarantees of [wang2024convergence] to hold (Conjecture 1). Nevertheless, we made substantial progress towards resolving this open question, which is also recurrent in the realizable process tomography literature [nadimpalli2024pauli, chen2023testing, bao2023testing]. Now, we summarize our results and the techniques used for each class.
@@
First, we consider any finite-sized class $\mathcal{C}$ of unitaries. Similarly to [zhao2023learning], we use Clifford classical shadows [huang2020predicting] to estimate the average infidelity between the unknown quantum channel $\Phi$ and all $U_{i}\in\mathcal{C}$. Then, we output the unitary in $\mathcal{C}$ with the smallest estimated distance. Thus, we solve proper agnostic unitary estimation for $\mathcal{C}$ with respect to the root mean squared trace distance $\mathrm{d_{avg}}$ and show this uses $\mathcal{O}(\log|\mathcal{C}|)$ queries to $\Phi$ and time $\mathcal{O}(|\mathcal{C}|)$. For many classes of interest, this results in a sample-efficient but computationally inefficient algorithm, emphasizing the importance of computational efficiency for agnostic learning. We show a weaker version of this result for agnostic process tomography of $\mathcal{C}$, which stems from converting from average infidelity to average trace distance for mixed inputs. As a corollary, by utilizing known covering number bounds [zhao2023learning], we obtain agnostic learning algorithms for the class of unitaries implementable by $G$ two-qubit gates. Note that in the realizable setting, [zhao2023learning] proves computational lower bounds. We prove this result in Sections 4.1 and 5.1.
@@
Recently, agnostic state tomography for several concept classes with respect to fidelity [grewal2024agnostic, chen2024stabilizer] has been studied. We show that this translates into guarantees for agnostic unitary estimation/process tomography for a compatible concept class $\mathcal{C}$ of unitaries with respect to $d_{F}$ with a quadratic increase in the error $\epsilon$. In particular, we require that the Choi state of each $U\in\mathcal{C}$ is in the concept class for agnostic state tomography. Our algorithm is simple: apply the agnostic state tomography algorithm to the Choi state of the unknown quantum process. Specifically, we obtain improper $1$-agnostic unitary estimation but a weaker error bound for agnostic process tomography (see Table 1(b)), again due to converting between infidelity and trace distance for mixed inputs. The key step in our proof shows a relationship between the infidelity and the normalized Frobenius distance between two Choi states. We prove this result in Sections 4.2 and 5.2.
@@
Next, we consider the class of Pauli strings $\{I,X,Y,Z\}^{\otimes n}$. We use the fact that one can efficiently sample from the distribution formed by the Pauli/Fourier coefficients [montanaro2008quantum]. After collecting samples, we simply output the most frequently sampled Pauli string. Thus, we achieve proper 1-agnostic unitary estimation/process tomography. We prove this in Sections 4.3 and 5.3.
@@
Pauli channels are a naturally occurring and widely studied [erhard2019characterizing, carignan2023error, harper2020efficient, tuckett2018ultrahigh, van2023probabilistic, ferracin2024efficiently, kim2023evidence, harper2020efficient, flammia2020efficient, chen2022quantum, flammia2021pauli, fawzi2023lower, chen2024tight, harper2021fast] class of quantum channels. Pauli channels are of the form $\sum_{x}p_{x}\sigma_{x}(\cdot)\sigma_{x}$ for a probability distribution $\{p_{x}\}_{x}$. Similarly to Pauli strings, we can sample from $\{p_{x}\}_{x}$. Then, agnostic tomography of Pauli channels reduces to empirically estimating $\{p_{x}\}_{x}$. Thus, we obtain proper 1-agnostic process tomography of Pauli channels. We prove this in Section 5.3.
@@
A $k$-junta unitary/channel is an $n$-qubit unitary/channel that acts only on $k$ qubits (Definitions 15 and 18). [chen2023testing, bao2023testing] devised realizable algorithms for learning junta unitaries/channels. Here, they estimate the influence (Equation 2.16) of each qubit, then select a highly influential subset of qubits and perform full state tomography on the Choi state restricted to these qubits.
@@
[arunachalam2024learning] develops realizable algorithms leveraging Bohnenblust-Hille (BH) inequalities for operators [volberg2024noncommutative]/
superoperators [arunachalam2024learning]. These inequalities provide upper bounds on norms of the spectra of low-degree objects. Using these bounds, [arunachalam2024learning] show that it suffices to learn significantly fewer Pauli/Fourier coefficients, resulting in algorithms with complexities independent of $n$. Using similar arguments for agnostic unitary estimation, we show that the algorithm of [arunachalam2024learning] is inefficient and only performs better than the trivial algorithm when the optimal error of the class of low-degree unitaries is at most inverse exponential in $n$, i.e., when the problem is almost realizable. We remark that improved bounds for the BH inequality may result in better guarantees for this algorithm and leave this open for future work. As this algorithm does not work well in the simpler unitary case, we do not analyze its performance for agnostic process tomography. We prove these results in Sections 4.5 and 5.5.
@@
The circuit class $\mathsf{QAC}^{0}$ consists of constant-depth circuits with controlled-$Z$ gates and arbitrary single-qubit gates [moore1999quantum]. In [nadimpalli2024pauli], the authors study the class of channels formed by applying a $\mathsf{QAC}^{0}$ circuit and tracing out all but one qubit. [nadimpalli2024pauli] shows that the Choi representations of such channels have Pauli weight concentrated on low-degree terms. Similarly to [nadimpalli2024pauli], we learn the low-degree coefficients of the Choi representation of an arbitrary channel using classical shadows [huang2020predicting], resulting in improper 1-agnostic process tomography for this class of $\mathsf{QAC}^{0}$ channels. We show this in Section 5.6.